3b+2c=46
-5c+b=-11

Answer by checkley75(3666)

-5c+b=-11
b=5c-11 now substitute (5c-11) for b in the other equation & solve for b
3(5c-11)+2c=46
15c-33+2c=46
17c=46+33
17c=79
c=79/17 answer
now substitute 79/17 for c and solve for b
b=5*79/17-11
b=395/17-11
b=(395-187)/17
b=208/17 answer
proof
-5*79/17+208/17=-11
(-395+208)/17=-11
-187/17=-11
-11=-11